OpenStudy (sbrummitt):

Solve for x. 5x - 1 = 26 plezz help

OpenStudy (wolfman77):

Same as the other one we did. Simplifying 5x + -1 = 26 Reorder the terms: -1 + 5x = 26 Solving -1 + 5x = 26 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '1' to each side of the equation. -1 + 1 + 5x = 26 + 1 Combine like terms: -1 + 1 = 0 0 + 5x = 26 + 1 5x = 26 + 1 Combine like terms: 26 + 1 = 27 5x = 27 Divide each side by '5'. x = 5.4 Simplifying x = 5.4 Make sense?

OpenStudy (wolfman77):

Do you understand this?

OpenStudy (sbrummitt):

yes thank you for the help.

OpenStudy (wolfman77):

Yup. No Problem! :D